OpenStudy (anonymous):

if alpha , beta r the o's of 2x^2-5x+7. find a polynomial those o's r 2alpha+3beta n 3alpha+2beta

OpenStudy (anonymous):

fo you know how to find the zeroes of a quadratic expression?

OpenStudy (anonymous):

let the polynomial be ax^2 +bx +c

OpenStudy (anonymous):

you know that alpha + beta = 5/2 and alpha times beta = 7/2

OpenStudy (anonymous):

therefore the conditions on the coeffiecents of the new polyminal are -b/a = 5 ( alpha +beta ) so -b/a = 25/2

OpenStudy (anonymous):

and the product of the two roots = 6alpha^2 +13aplha beta +6beta^2 = c/a

OpenStudy (anonymous):

just easy simultaneous eqns u do the rest